2) A student measured the magnetic susceptibility of a transition metal complex using the same MSB instrument used in this laboratory experiment. A mass of 0.163 grams of the compound with a molar mass of 445 g/mol was transferred into an MSB tube to produce a sample length (L) of 1.74 cm. The net magnetic susceptibility reading (R) of 2304.0 nerg·G–2·cm–1 was measured at 25°C.

a) Calculate the molar magnetic susceptibility, χm, in erg·G–2·mol–1.

b) Calculate the magnetic moment-Bohr magneton ratio, μ'.

c) Estimate the number of unpaired electrons in the complex.
Answer #1

a) The net magnetic susceptibility reading (R) of 2304.0 nerg·G–2·cm–1 was measured at 25°C

Number of mole of compound (n)= 0.163/445=3.663 x 10-4

so molar the molar magnetic susceptibility, χm, = n x net magnetic susceptibility x a sample length =

2304.0 nerg·G–2·cm–1 x  1.74 cm / 3.663 x 10-4 mol = 1.095*10-2erg·G–2·mol–1.

b)the magnetic moment-Bohr magneton ratio is given by

\mu =2.84\sqrt{\chi T} ; 2.84\sqrt{0.01095 *298} = 2.84 *1.806= 5.12

C) µs = g \sqrt{S (S+1)} ; so nearest value of S = 4 since \sqrt{4 *5} .

so Number of unpaired electron are 4.

Once this is ratio is calculated, then we can calculate